When John's age 2yrs ago was added to his age 5yrs to come the sum was 35 how old is John now?
Answer by ewatrrr(24785)

x John's age Now
x-2 + x+5 = 35
2x + 3 = 35
2x = 32
x = 16
And...checking
14 + 21 = 35